# Distribution of local signs of modular forms and murmurations of Fourier coefficients

### v2

@@ 3.1.1. Computing $\tilde{A}_{1,\varepsilon}(r;M)$ when $4\ell<q^{r+2\varepsilon}$
In the following analysis, we will assume $\ell$ is squarefree. Write $-4q^{r}\ell=\lambda^{2}\Delta_{0}$, where $\Delta_{0}$ is a fundamental discriminant. We can rewrite $H(-4q^{r}\ell)$ and $H(-q^{r}\ell)$ in terms of $h^{\prime}(\Delta_{0})$ on a case-by-case basis as in Table 1. We will use this to calculate $\tilde{A}_{1}$ in cases.
The following computation will be useful. First, for an integer $\Delta$, note that (${\Delta\overwithdelims( )\boldsymbol{\cdot}}*|\mu|$ is the multiplicative function given on nontrivial prime powers by
($({\Delta\overwithdelims( )\boldsymbol{\cdot}}*|\mu|)(p^{m})=\begin{cases}1+{\Delta\overwithdelims( )p}&\text{if }p\nmid\Delta\text{ or }m=1,\\ 0&\text{if }p\mid\Delta\text{ and }m=1.\end{cases}$
@@
(3.5) $\kappa_{\Delta}(p^{m})=\begin{cases}{\Delta\overwithdelims( )p}-1&\text{if }m=1,\\ -{\Delta\overwithdelims( )p}&\text{if }p\nmid\Delta\text{ and }m=2,\\ -1&\text{if }p\mid\Delta\text{ and }m=2,\\ 1&\text{if }p\mid\Delta\text{ and }m=3,\\ 0&\text{else}.\end{cases}$
@@
$r$ even, $\ell\not\equiv 3\bmod 4$ $-4\ell$ $q^{r/2}$ $\eta_{-4\ell}(q^{r/2})h^{\prime}(-4\ell)$ 00
$r$ even, $\ell\equiv 3\bmod 4$$\ell$ $2q^{r/2}$ $\eta_{-\ell}(2q^{r/2})h^{\prime}(-\ell)$ $\eta_{-\ell}(q^{r/2})h^{\prime}(-\ell)$
$r$ odd, $q\ell\not\equiv 3\bmod 4$ $-4q\ell$ $q^{(r-1)/2}$ $\eta_{-4q\ell}(q^{(r-1)/2})h^{\prime}(-4q\ell)$ 00
$r$ odd, $q\ell\equiv 3\bmod 4$ $-q\ell$ $2q^{(r-1)/2}$ $\eta_{-q\ell}(2q^{(r-1)/2})h^{\prime}(-q\ell)$ $\eta_{-q\ell}(q^{(r-1)/2})h^{\prime}(-q\ell)$
@@
where we interpret $H_{t}$ to be 00 if $t\not\in\mathbb{Z}$. Hence
@@ 4. Dimension formulas
In particular $\kappa_{-q^{r}}(M)=0$ if and only if (i) (${-q^{r}\overwithdelims( )p}=1$ for some $p\parallel M$; or (ii) $v_{p}(M)\geq 3$ for some $p$. Assuming neither (i) nor (ii) hold, then $\kappa_{-q^{r}}(M)=(-1)^{\omega_{2}(-q^{r};M)}(-2)^{\omega_{1}(M)}$, where $\omega_{1}(M)$ is the number of primes sharply dividing $M$ and $\omega_{2}(n;M)$ is the number of $p^{2}\parallel M$ such that (${n\overwithdelims( )p}=1$.
We also tabulate the values of $\alpha_{1}(-q^{r};e)$ by cases in Table 2. These calculations use the fact that $H(-4q^{r})=(3-{-q\overwithdelims( )2})H(-q^{r})$ when $-q^{r}\equiv 1\bmod 4$. In particular, one sees that for $q$ odd, $\alpha_{1}(-q^{r};e)=0$ if and only if (i) $q^{r}\equiv 1\bmod 4$ and $e\geq 4$; (ii) $q\equiv 3\bmod 8$, $r$ is odd and $e\geq 5$; or (iii) $q\equiv 7\bmod 8$, $r$ is odd and $e>0$. We note that $\alpha_{1}(-q^{r};e)\leq 0$ if (a) $e\neq 1,2$ or (b) $e=3$, $r$ is odd, and $q\equiv 7\bmod 8$. Otherwise $\alpha_{1}(-q^{r};e)\geq 0$.
@@ 00 $H(-4q^{r})$ $(3-{-q\overwithdelims( )2})H(-q^{r})$ $H(-2^{r+2})$
$1,2$ $-H(-4q^{r})$ ($\left({-q\overwithdelims( )2}-1\right)H(-q^{r})$
$3$ $H(-4q^{r})$ $3\left({-q\overwithdelims( )2}-1\right)H(-q^{r})$
$4$ 00 $2\left(1-2{-q\overwithdelims( )2}\right)H(-q^{r})$
$\geq 5$ 00 00
@@ 4.1. Dimensions for $r=1$
    Suppose $q\geq 5$ and either $k\geq 4$ or $M$ is not squarefree. Then $\Delta_{k}(q,M)=0$ if and only if (i) $M^{\prime}$ is not cubefree; (ii) (${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$; (iii) $16\mid M$ and $q\equiv 1\bmod 4$; (iv) $32\mid M$ and $q\equiv 3\bmod 8$; or (v) $v_{2}(M)\neq 0,4$ and $q\equiv 7\bmod 8$.
@@
    Suppose $q=2$. We have $\Delta_{k}(2,M)=0$ if and only if (i) $M$ is not cubefree; (ii) $\prod_{p^{2}\parallel M}{2\overwithdelims( )p}$ is $-1$ if $k\equiv 0,2\bmod 8$ and $+1$ if $k\equiv 4,6\bmod 8$; or (iii) $k=2$ and $M=1$ or $M\equiv 3,5\bmod 8$ is prime.
@@
    Suppose $q=3$. If $k\geq 4$ or $M$ is not squarefree, then $\Delta_{k}(3,M)=0$ if and only if (i) $M^{\prime}$ is not cubefree; (ii) (${-3\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$; (iii) $32\mid M$; (iv) $M$ is odd and $k\equiv 4,10\bmod 12$; or (v) $4\parallel M$ and $k\not\equiv 4,10\bmod 12$. If $k=2$ and $M$ is squarefree, then $\dim S_{k}^{\mathrm{new}}(3M)^{+_{3}}=\dim S_{k}^{\mathrm{new}}(3M)^{-_{3}}$ if and only if $M=1,2$.
@@
When $M=1$, $\Delta_{k}(q,M)=0$ if and only if $H(-4q)=2$, i.e., if $q\in\{5,7,13,17\}$. When $M=2$, $\Delta_{k}(q,M)=0$ if and only if $H(-4q)=2H(-q)+2$, i.e., if and only if $q\in\{5,11,13,19,37,43,67,163\}$. When $M=p\geq 3$, $\Delta_{k}(q,M)=0$ if and only if $H(-4q)=1$ and (${-q\overwithdelims( )p}=-1$, which never happens. When $M=2p$, for a prime $p\geq 3$, $\Delta_{k}(q,M)=0$ if and only if $12H(-q)=H(-4q)+$ 1, which also never happens. This finishes case (2).
Case (3): Next suppose $q=2$, and $k\neq 2$ or $M$ is not squarefree. Then $t^{\mathrm{new}}(1;M)=0$ if and only if $\kappa_{-2}(M)=\kappa_{-1}(M)=0$ or $(-1)^{k/2}p_{k}(\sqrt{2},1)=\prod_{p^{2}\parallel M}{2\overwithdelims( )p}$. The former never happens. The latter condition means $\prod_{p^{2}\parallel M}{2\overwithdelims( )p}$ is $-1$ if $k\equiv 0,2\bmod 8$ and $+1$ if $k\equiv 4,6\bmod 8$.
@@
Hence if $k\neq 2$ or $M$ is not squarefree, then $\Delta_{k}(q,M)=0$ if and only if one of (i)–(iii) above holds; (iv) $M^{\prime}$ is not cubefree; or (v) (${-3\overwithdelims( )p}=1$ for some odd $p\parallel M$.
@@ 4.2. Dimensions for $r\geq 3$ odd
Recall $\kappa_{-q}(M^{\prime})=0$ if and only if $M^{\prime}$ is not cubefree or (${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$.
@@
where $\Delta_{0}=-4q$ if $q\equiv 1,2\bmod 4$ and $\Delta_{0}=-q$ if $q\equiv 3\bmod 4$. Here $c_{1}\in\{0,\pm 1,\pm 2,4,-6\}$ according to the cases in Table 2 (and (${-q\overwithdelims( )2}$ when $q\equiv 3\bmod 4$). In particular, $c_{1}=0$ if and only if (a) $e\geq 5$, (b) $e=4$ and $q\equiv 1\bmod 4$, or (c) $e=1,2,3$ and $q\equiv 7\bmod 8$.
Moreover, by Table 1 and (2.2), we see that
$\aleph$ $=c_{1}\left(\eta_{\Delta_{0}}(q^{\frac{r-1}{2}})-\eta_{\Delta_{0}}(2q^{\frac{r-3}{2}})+\delta_{r\geq 5}\eta_{\Delta_{0}}(q^{\frac{r-5}{2}})\right)h^{\prime}(\Delta_{0})$
$=c_{1}\left(\sigma(q^{\frac{r-1}{2}})-\sigma(2q^{\frac{r-3}{2}})+\delta_{r\geq 5}\sigma(q^{\frac{r-5}{2}})\right)h^{\prime}(\Delta_{0}).$
@@
Let $q$ be a prime, $r\geq 3$ odd, $M\geq 1$ be coprime to $q$, and write $M=2^{e}M^{\prime}$, where $M^{\prime}$ is odd. Assume $q^{r}\neq 8,27$. Then $\Delta_{k}(q^{r},M)=0$ if and only if (i) $M^{\prime}$ is not cubefree; (ii) (${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$; (iii) $e\geq 5$; (iv) $e=4$ and $q\equiv 1\bmod 4$; or (v) $e=1,2,3$ and $q\equiv 7\bmod 8$.
@@ 4.3. Dimensions for $r=2$
Recall that $\tilde{A}_{2}(r;M)=0$ unless $M$ is a square and $q^{r}\in\{1,4\}$, in which case it is $-\frac{1}{2}\alpha_{2}(M)$. Note that $0\leq\alpha_{2}(M)<\sqrt{M}$. This combined with our analysis below will imply that the $\tilde{A}_{2}$ terms in $\Delta_{k}(q^{2},M)$ will not contribute to the main asymptotics in either case.
@@ 4.4. Dimensions for $r\geq 4$ even
=$\aleph=\begin{cases}2^{\frac{r-4}{2}}&\text{if }q=2,\\ \frac{1}{2}q^{\frac{r-4}{2}}(q-1)(q-{-1\overwithdelims( )q})&\text{if }q\neq 2\text{ and }e=0,3,\\ -\frac{1}{2}q^{\frac{r-4}{2}}(q-1)(q-{-1\overwithdelims( )q})&\text{if }q\neq 2\text{ and }e=1,2,\\ 0&\text{if }e\geq 4.\end{cases}$
@@ 5.1. Traces for small $\ell$
Assume $\ell<\frac{q}{4}$ is prime. Suppose either $M^{\prime}$ is not cubefree or $32\mid M$. Then the trace of $T_{\ell}$ on $S_{k}^{\mathrm{new}}(qM)^{\pm_{q}}$ is independent of the sign ±$\pm_{q}$ for $\ell<\frac{q}{4}$. Equivalently, since $\Delta_{k}(q,M)=0$, the average of $a_{\ell}(f)$ over newforms in $S_{k}^{\mathrm{new}}(qM)^{+_{q}}$ is equal to that for $S_{k}^{\mathrm{new}}(qM)^{-_{q}}$ for primes $\ell<\frac{q}{4}$.
@@
Suppose $M$ is cubefree and $\ell<\frac{q}{4}$ is prime. If either (i) (${-q\ell\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$ or (ii) $e=1,2$ and $q\ell\equiv 7\bmod 8$, then
@@
Now we compare the sign of this $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}W_{q}$ (or whether it is 0 or not) with the sign of $\Delta_{k}(q,M)$. For simplicity, assume $k\geq 4$ or $M$ is not squarefree so that the $\delta_{k=2}\mu(M)\sigma(\ell)$ term vanishes. Then from 4.1, we have $\Delta_{k}(q,M)=0$ if and only if (i) (${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$ or (ii) $M$ is even and $q\equiv 7\bmod 8$.
@@
for $q$ large $H(-4q\ell)$ dominates $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}$. Thus the sign of $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)^{\pm q}}T_{\ell}$ will be ±$\pm 1$ times the sign of $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}W_{q}$ for large $q$ such that the latter trace is nonzero.

### v3

@@ 3.1.1. Computing $\tilde{A}_{1,\varepsilon}(r;M)$ when $4\ell<q^{r+2\varepsilon}$
In the following analysis, we will assume $\ell$ is squarefree. Write $4q^{r}\ell=\lambda^{2}\Delta_{0}$, where $\Delta_{0}$ is a fundamental discriminant. We can rewrite $H(-4q^{r}\ell)$ and $H(-q^{r}\ell)$ in terms of $h^{\prime}(\Delta_{0})$ on a case-by-case basis as in Table 1. We will use this to calculate $\tilde{A}_{1}$ in cases.
The following computation will be useful. First, for an integer $\Delta$, note that ${\Delta\overwithdelims( )\boldsymbol{\cdot}}*|\mu|$ is the multiplicative function given on nontrivial prime powers by
$({\Delta\overwithdelims( )\boldsymbol{\cdot}}*|\mu|)(p^{m})=\begin{cases}1+{\Delta\overwithdelims( )p}&\text{if }p\nmid\Delta\text{ or }m=1,\\ 0&\text{if }p\mid\Delta\text{ and }m=1.\end{cases}$
@@
[…]
@@
$r$ even, $\ell\not\equiv 3\bmod 4$ $4\ell$ $q^{r/2}$ $\eta_{-4\ell}(q^{r/2})h^{\prime}(-4\ell)$ 0
$r$ even, $\ell\equiv 3\bmod 4$ -$\ell$ $2q^{r/2}$ $\eta_{-\ell}(2q^{r/2})h^{\prime}(-\ell)$ $\eta_{-\ell}(q^{r/2})h^{\prime}(-\ell)$
$r$ odd, $q\ell\not\equiv 3\bmod 4$ $4q\ell$ $q^{(r-1)/2}$ $\eta_{-4q\ell}(q^{(r-1)/2})h^{\prime}(-4q\ell)$ 0
$r$ odd, $\ell\equiv 3\bmod 4 q-q\ell$ $2q^{(r-1)/2}$ $\eta_{-q\ell}(2q^{(r-1)/2})h^{\prime}(-q\ell)$ $\eta_{-q\ell}(q^{(r-1)/2})h^{\prime}(-q\ell)$
@@
where we interpret $H_{t}$ to be 0 if $t\not\in\mathbb{Z}$. Hence
@@ 4. Dimension formulas
In particular $\kappa_{-q^{r}}(M)=0$ if and only if (i) ${-q^{r}\overwithdelims( )p}=1$ for some $p\parallel M$; or (ii) $v_{p}(M)\geq 3$ for some $p$. Assuming neither (i) nor (ii) hold, then $\kappa_{-q^{r}}(M)=(-1)^{\omega_{2}(-q^{r};M)}(-2)^{\omega_{1}(M)}$, where $\omega_{1}(M)$ is the number of primes sharply dividing $M$ and $\omega_{2}(n;M)$ is the number of $p^{2}\parallel M$ such that ${n\overwithdelims( )p}=1$.
We also tabulate the values of $\alpha_{1}(-q^{r};e)$ by cases in Table 2. These calculations use the fact that $H(-4q^{r})=(3-{-q\overwithdelims( )2})H(-q^{r})$ when $q^{r}\equiv 1\bmod 4$. In particular, one sees that for $q$ odd, $\alpha_{1}(-q^{r};e)=0$ if and only if (i) $q^{r}\equiv 1\bmod 4$ and $e\geq 4$; (ii) $q\equiv 3\bmod 8$, $r$ is odd and $e\geq 5$; or (iii) $q\equiv 7\bmod 8$, $r$ is odd and $e>0$. We note that $\alpha_{1}(-q^{r};e)\leq 0$ if (a) $e\neq 1,2$ or (b) $e=3$, $r$ is odd, and $q\equiv 7\bmod 8$. Otherwise $\alpha_{1}(-q^{r};e)\geq 0$.
@@ 0 $H(-4q^{r})$ $(3-{-q\overwithdelims( )2})H(-q^{r})$ $H(-2^{r+2})$
$1,2$ $H(-4q^{r})$ $\left({-q\overwithdelims( )2}-1\right)H(-q^{r})$
33 $H(-4q^{r})$ $3\left({-q\overwithdelims( )2}-1\right)H(-q^{r})$
44 0 $2\left(1-2{-q\overwithdelims( )2}\right)H(-q^{r})$
$\geq 5$ 0 0
@@ 4.1. Dimensions for $r=1$
    Suppose $q\geq 5$ and either $k\geq 4$ or $M$ is not squarefree. Then $\Delta_{k}(q,M)=0$ if and only if (i) $M^{\prime}$ is not cubefree; (ii) ${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$; (iii) $16\mid M$ and $q\equiv 1\bmod 4$; (iv) $32\mid M$ and $q\equiv 3\bmod 8$; or (v) $v_{2}(M)\neq 0,4$ and $q\equiv 7\bmod 8$.
@@
    Suppose $q=2$. We have $\Delta_{k}(2,M)=0$ if and only if (i) $M$ is not cubefree; (ii) $\prod_{p^{2}\parallel M}{2\overwithdelims( )p}$ is 1-1 if $k\equiv 0,2\bmod 8$ and +1+1 if $k\equiv 4,6\bmod 8$; or (iii) $k=2$ and $M=1$ or $M\equiv 3,5\bmod 8$ is prime.
@@
    Suppose $q=3$. If $k\geq 4$ or $M$ is not squarefree, then $\Delta_{k}(3,M)=0$ if and only if (i) $M^{\prime}$ is not cubefree; (ii) ${-3\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$; (iii) $32\mid M$; (iv) $M$ is odd and $k\equiv 4,10\bmod 12$; or (v) $4\parallel M$ and $k\not\equiv 4,10\bmod 12$. If $k=2$ and $M$ is squarefree, then $\dim S_{k}^{\mathrm{new}}(3M)^{+_{3}}=\dim S_{k}^{\mathrm{new}}(3M)^{-_{3}}$ if and only if $M=1,2$.
@@
When $M=1$, $\Delta_{k}(q,M)=0$ if and only if H(4q)=2H(-4q)=2, i.e., if $q\in\{5,7,13,17\}$. When $M=2$, $\Delta_{k}(q,M)=0$ if and only if H(4q)=2H(q)+2H(-4q)=2H(-q)+2, i.e., if and only if $q\in\{5,11,13,19,37,43,67,163\}$. When $M=p\geq 3$, $\Delta_{k}(q,M)=0$ if and only if H(4q)=1H(-4q)=1 and ${-q\overwithdelims( )p}=-1$, which never happens. When $M=2p$, for a prime $p\geq 3$, $\Delta_{k}(q,M)=0$ if and only if 2H(q)=H(4q)+12H(-q)=H(-4q)+1, which also never happens. This finishes case (2).
Case (3): Next suppose $q=2$, and $k\neq 2$ or $M$ is not squarefree. Then $t^{\mathrm{new}}(1;M)=0$ if and only if $\kappa_{-2}(M)=\kappa_{-1}(M)=0$ or $(-1)^{k/2}p_{k}(\sqrt{2},1)=\prod_{p^{2}\parallel M}{2\overwithdelims( )p}$. The former never happens. The latter condition means $\prod_{p^{2}\parallel M}{2\overwithdelims( )p}$ is 1-1 if $k\equiv 0,2\bmod 8$ and +1+1 if $k\equiv 4,6\bmod 8$.
@@
Hence if $k\neq 2$ or $M$ is not squarefree, then $\Delta_{k}(q,M)=0$ if and only if one of (i)–(iii) above holds; (iv) $M^{\prime}$ is not cubefree; or (v) ${-3\overwithdelims( )p}=1$ for some odd $p\parallel M$.
@@ 4.2. Dimensions for $r\geq 3$ odd
Recall $\kappa_{-q}(M^{\prime})=0$ if and only if $M^{\prime}$ is not cubefree or ${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$.
@@
where $\Delta_{0}=-4q$ if $q\equiv 1,2\bmod 4$ and $\Delta_{0}=-q$ if $q\equiv 3\bmod 4$. Here $c_{1}\in\{0,\pm 1,\pm 2,4,-6\}$ according to the cases in Table 2 (and $q\overwithdelims( )2$} when $q\equiv 3\bmod 4$). In particular, $c_{1}=0$ if and only if (a) $e\geq 5$, (b) $e=4$ and $q\equiv 1\bmod 4$, or (c) $e=1,2,3$ and $q\equiv 7\bmod 8$.
Moreover, by Table 1 and (2.2), we see that [^4] In the published version, the factor of 2 mistakenly appeared inside the arguments of $\eta_{\Delta_{0}}$ and $\sigma$. This final simplified expression for $\aleph$ did not appear in the published version.
$\aleph$ $=c_{1}\left(\eta_{\Delta_{0}}(q^{\frac{r-1}{2}})-2\eta_{\Delta_{0}}(q^{\frac{r-3}{2}})+\delta_{r\geq 5}\eta_{\Delta_{0}}(q^{\frac{r-5}{2}})\right)h^{\prime}(\Delta_{0})$
$=c_{1}\left(\sigma(q^{\frac{r-1}{2}})-2\sigma(q^{\frac{r-3}{2}})+\delta_{r\geq 5}\sigma(q^{\frac{r-5}{2}})\right)h^{\prime}(\Delta_{0})$
$=c_{1}\left(q^{\frac{r-1}{2}}-q^{\frac{r-3}{2}}\right)h^{\prime}(\Delta_{0}).$
@@
Let $q$ be a prime, $r\geq 3$ odd, $M\geq 1$ be coprime to $q$, and write $M=2^{e}M^{\prime}$, where $M^{\prime}$ is odd. Assume $q^{r}\neq 8,27$. Then $\Delta_{k}(q^{r},M)=0$ if and only if (i) $M^{\prime}$ is not cubefree; (ii) ${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$; (iii) $e\geq 5$; (iv) $e=4$ and $q\equiv 1\bmod 4$; or (v) $e=1,2,3$ and $q\equiv 7\bmod 8$.
@@ 4.3. Dimensions for $r=2$
Recall that $\tilde{A}_{2}(r;M)=0$ unless $M$ is a square and $q^{r}\in\{1,4\}$, in which case it is $\frac{1}{2}\alpha_{2}(M)$. Note that $0\leq\alpha_{2}(M)<\sqrt{M}$. This combined with our analysis below will imply that the $\tilde{A}_{2}$ terms in $\Delta_{k}(q^{2},M)$ will not contribute to the main asymptotics in either case.
@@ 4.4. Dimensions for $r\geq 4$ even
[…]
@@ 5.1. Traces for small $\ell$
Assume $\ell<\frac{q}{4}$ is prime. Suppose either $M^{\prime}$ is not cubefree or $32\mid M$. Then the trace of $T_{\ell}$ on $S_{k}^{\mathrm{new}}(qM)^{\pm_{q}}$ is independent of the sign $\pm_{q}$ for $\ell<\frac{q}{4}$. Equivalently, since $\Delta_{k}(q,M)=0$, the average of $a_{\ell}(f)$ over newforms in $S_{k}^{\mathrm{new}}(qM)^{+_{q}}$ is equal to that for $S_{k}^{\mathrm{new}}(qM)^{-_{q}}$ for primes $\ell<\frac{q}{4}$.
@@
Suppose $M$ is cubefree and $\ell<\frac{q}{4}$ is prime. If either (i) ${-q\ell\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$ or (ii) $e=1,2$ and $q\ell\equiv 7\bmod 8$, then
@@
Now we compare the sign of this $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}W_{q}$ (or whether it is 0 or not) with the sign of $\Delta_{k}(q,M)$. For simplicity, assume $k\geq 4$ or $M$ is not squarefree so that the $\delta_{k=2}\mu(M)\sigma(\ell)$ term vanishes. Then from 4.1, we have $\Delta_{k}(q,M)=0$ if and only if (i) ${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$ or (ii) $M$ is even and $q\equiv 7\bmod 8$.
@@
for $q$ large $H(-4q\ell)$ dominates $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}$. Thus the sign of $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)^{\pm q}}T_{\ell}$ will be $\pm 1$ times the sign of $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}W_{q}$ for large $q$ such that the latter trace is nonzero.